OpenStudy (anonymous):

help!!!! The function fx=(1.006235)^12x models the monthly interest that a bank offers to dan after x years. Dan converts the function to have x isolated in the exponent. What is the approximate rate of growth?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm f(x)=(1.006235)^{12x}\] An important rule of exponents to remember:\[\Large\rm \color{royalblue}{a^{bc}=(a^b)^c}\]

zepdrix (zepdrix):

Applying the rule allows us to isolate our x: \[\Large\rm f(x)=\left(1.006235^{12}\right)^x\]

OpenStudy (anonymous):

so I can do 1.006235^12 ?

zepdrix (zepdrix):

Yah, what does that give you? :)

OpenStudy (anonymous):

1.077439845599 so 1%?

zepdrix (zepdrix):

the 1. represents 100%. You can split it up like this if it will help to see what's going on.\[\Large\rm (1.0774)^x=(1 +\color{orangered}{.0774})^x\]The 1 represents the previous whole amount. The orange part is the rate of grown portion.

OpenStudy (anonymous):

so its A?

zepdrix (zepdrix):

So what is that value, that decimal, written as a percent?\[\Large\rm .0774 =\text{_%}\]

zepdrix (zepdrix):

No, it is not A.

OpenStudy (anonymous):

7.74

OpenStudy (anonymous):

WHICH 8%

zepdrix (zepdrix):

Approximately 8%? Yayyy good job \c:/

OpenStudy (anonymous):

THAT MAKES SENSE!! LOL THANKS :)

OpenStudy (anonymous):

CAN YOU HELP ME ON SIMPLIFYING?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

CAN YOU HELP ME WITH #6? I TOTALLY FORGOT HOW TO DO IT

zepdrix (zepdrix):

Rationalize your root term using this rule:\[\Huge\rm \color{royalblue}{\sqrt[a]{x}}=x^{1/a}\]

OpenStudy (anonymous):

SO 8^1/5?

zepdrix (zepdrix):

Yes, good. Let's do it to the top AND bottom ok? \[\Large\rm \frac{\sqrt[2]{8}}{\sqrt[5]{8}}=\frac{8^{1/2}}{8^{1/5}}\]

OpenStudy (anonymous):

WHERE DID YOU GET THE 2 FROM? THERE IS NO 2 IN THE QUESTION

zepdrix (zepdrix):

When there is no index written on the root, it is implied to the `square` root, second root.

OpenStudy (anonymous):

OH OKAY

zepdrix (zepdrix):

implied to be the*

zepdrix (zepdrix):

Then our rules of exponents tells us,\[\Large\rm \color{orangered}{\frac{x^a}{x^b}=x^{a-b}}\]

OpenStudy (anonymous):

5/10-2/10?

OpenStudy (anonymous):

RIGHT?

zepdrix (zepdrix):

Mmm yes, good!

OpenStudy (anonymous):

IM SORRY BUT CAN YOU TELL ME IF I GOT THIS ONE RIGHT?

zepdrix (zepdrix):

Uh ohhhh I think you made a boo boo.

zepdrix (zepdrix):

You've got the right idea, since we're shifting to the `left`, we `add` to our x, yah? But we're adding 13 to our x+1.

OpenStudy (anonymous):

DANG! oh yeah I forgot about that, so its B?

zepdrix (zepdrix):

because of the + 1? yay good job \c:/

OpenStudy (anonymous):

OKAY COOL :)

OpenStudy (anonymous):

im working on a problem now once I formulate my answer can you check it for me?

OpenStudy (anonymous):

@zepdrix okay, I got the numerator but I'm not sure about the denominator.

zepdrix (zepdrix):

You got the numerator figured out?\[\Large\rm \sqrt[5]{8}\sqrt{8}=8^{1/5}\cdot 8^{1/2}=8^{1/5+1/2}=8^{7/10}\]That what you came up with for your numerator?

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

but what about the denominator?

zepdrix (zepdrix):

Ohhh ok, lemme state the rule from earlier but with a power introduced.\[\Huge\rm \color{royalblue}{\sqrt[a]{x^b}}=x^{b/a}\]

OpenStudy (anonymous):

is it c? 8^-29/30

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

thanks, can we do another problem?

OpenStudy (anonymous):

only like two more pretttty please?

zepdrix (zepdrix):

you... -_-

OpenStudy (anonymous):

no? lol okay I understand

OpenStudy (anonymous):

i just want to make sure m doing it correctly

zepdrix (zepdrix):

you..... you and your ways... -_- Yah it's fine if I'm around c: Just toss a @ if I'm not responding or whatev

OpenStudy (anonymous):

okay thanks so much !!

OpenStudy (anonymous):

its either B OR D, but im going with B

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

NO D?

zepdrix (zepdrix):

Do you remember the very first problem we did? \[\Large\rm (1 +\color{orangered}{.0774})^x\]The orange represented the MORE each month. The 1 + orange represents `the full amount + some more each month`

zepdrix (zepdrix):

So if we have an increase of 5% each month, next month she'll have 105% of her pop music from last month, yes?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

AND IN ORDER TO GET THAT MOVE CIMAL PLACE OVER BETWEEN THE 1 AND THE 0

OpenStudy (anonymous):

BUT WAS I CORRECT?

zepdrix (zepdrix):

lol those caps, simmer down :o

OpenStudy (anonymous):

lol sorry, they don't mean anything

zepdrix (zepdrix):

Which one did you go with? Yah you need to move the decimal over to get the correct value, 1.05. Just remember that since it's an `increase` we want the original amount + the 5%. 105%

OpenStudy (anonymous):

D?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

OKAY! last one. this one i'm not sure how to do at all

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Oh yah this question is quite a bit harder than the previous ones. So to find `rate of change` or `slope` we use the slope formula:

zepdrix (zepdrix):

So in this case, we want to know how the function changed from year 1 to year 5.

zepdrix (zepdrix):

\[\Large\rm \text{Change in students}=\frac{y_2-y_1}{x_2-x_1}\]Does that formula look familiar?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

We want to write it in function notation, which looks a little scarier.\[\Large\rm \frac{f(x_2)-f(x_1)}{x_2-x_1}\]And for this problem we're using this data: \(\Large\rm x_1=1, \text{ first year}\) \(\Large\rm x_2=5, \text{ fifth year}\)

zepdrix (zepdrix):

\[\Large\rm \text{Change in students}=\frac{f(5)-f(1)}{5-1}\]

OpenStudy (anonymous):

so is it just 1?

zepdrix (zepdrix):

No. f(5)-f(1) does not equal 4. We have to calculate the function at these values.\[\Large\rm f(\color{orangered}{x})=11(1.35)^{\color{orangered}{x}}\]Evaluating the function at 5 gives us:\[\Large\rm f(\color{orangered}{5})=11(1.35)^{\color{orangered}{5}}\]Which simplifies to,\[\Large\rm f(\color{orangered}{5})\approx 49.3\]

zepdrix (zepdrix):

Understand what I did to find f(5)? Think you can figure out the f(1)?

OpenStudy (anonymous):

yes, 14.85?

zepdrix (zepdrix):

Ok great, so our rate of change we can write as,\[\Large\rm \text{Change in students}\approx \frac{49.32-14.85}{5-1}\]

OpenStudy (anonymous):

8.6125?

zepdrix (zepdrix):

Mmmmm yah I think we did that correctly. That sounds about right :D

zepdrix (zepdrix):

Remember, we're talking about students so we don't want to consider a `part of a student`. We need to round to the nearest whole number.

OpenStudy (anonymous):

so 9

zepdrix (zepdrix):

cool c:

OpenStudy (anonymous):

thank you so much !!!!!!!!!!

zepdrix (zepdrix):

You might want to uhhh.. be more careful when you post your questions. I think you can get in trouble for posting exam questions. Either banned from the site and/or reported to flvs. So either edit your pictures so it only includes the question, not the exam details at the top of the page and all that. Or try to post questions from a practice test or something.

zepdrix (zepdrix):

I don't really care that much, but I don't wanna see you get in twouble :O

OpenStudy (anonymous):

okay thank you